### v4

# FREE: The Foundational Semantic Recognition for Modeling Environmental Ecosystems

Abstract
Modeling environmental ecosystems is critical for the sustainability of our planet, but is extremely challenging due to the complex underlying processes driven by interactions amongst a large number of physical variables. As many variables are difficult to measure at large scales, existing works often utilize a combination of observable features and locally available measurements or modeled values as input to build models for a specific study region and time period. This raises a fundamental question in advancing the modeling of environmental ecosystems: how to build a general framework for modeling the complex relationships amongst various environmental data over space and time? In this paper, we introduce a framework, FREE, which maps available environmental data into a text space and then converts the traditional predictive modeling task in environmental science to a semantic recognition problem. The proposed framework leverages recent advances in Large Language Models (LLMs) to supplement the original input features with natural language descriptions. This framework facilitates capturing the data semantics and allows harnessing the irregularities of input features. When used for long-term prediction, FREE has the flexibility to incorporate newly collected observations to enhance future prediction. The efficacy of FREE is evaluated in the context of two societally important real-world applications, predicting stream water temperature in the Delaware River Basin and predicting annual corn yield in Illinois and Iowa. Beyond the superior predictive performance over multiple baselines, FREE is shown to be more data- and computation-efficient as it can be pre-trained on simulated data generated by physics-based models.
1 Introduction
Understanding the dynamics of environmental ecosystems is critical for the sustainable management of natural resources and mitigating natural disasters such as algal blooms and floods. This is becoming even more important given the compelling need for food and water supply from a growing world population and a more unpredictable climate Mbow et al. (2020); Konar et al. (2011). Modeling environmental ecosystems is challenging as these systems are shaped by the complex interactions of a large number of physical variables, such as weather, soil, water, and plants. Hence, it often requires the combination of data from the network of weather stations, remote sensing, and field measurements to jointly model complex system dynamics. However, many physical variables cannot be easily measured over a large scale. They are often only available at certain locations or during certain time periods, or they are derived from models based on established physical theories. Even for those observable variables, they are often sparse due to the substantial cost required for the data collection (e.g., sensor deployment and field survey).
This paper focuses on two important applications, predicting stream water temperature and annual crop yield. Success in both of these prediction tasks can assist a range of applications in optimizing resource allocation and management strategies. For example, the changing climate is making it harder to grow crops in the US Corn Belt, threatening food supplies and farmer livelihoods; drinking water reservoir operators in the Delaware River Basin (our study region) need to supply safe drinking water to over 15 million people while also maintaining sufficient cool water temperatures in the river network downstream of the reservoirs Williamson et al. (2015). Given the importance of these two tasks in ensuring food and water security, scientists from multiple domains, including hydrology, agronomy, meteorology, and environmental engineering, have been developing physics-based models (also referred to as process-based models) to simulate underlying physical processes for different components of aquatic and agricultural ecosystems. Despite their extensive use, the majority of physics-based models are necessarily approximations of reality due to incomplete knowledge or excessive complexity in modeling certain processes. More recently, machine learning (ML) models have been increasingly considered as an alternative approach for modeling environmental ecosystems given their computational efficiency in the prediction phase and their ability to automatically extract complex data patterns. ML methods have found success in many environmental applications, including climate science Rolnick et al. (2022), hydrology Chen et al. (2022), and agriculture Liu et al. (2022).
However, current ML approaches face several major challenges in fully leveraging available data for modeling ecosystems. First, ML models often treat input physical variables as independent numerical features but do not explicitly capture the nature of these variables and their interdependent physical and ecological relationships. This is further exacerbated by the scarcity of observation data in many real-world environmental ecosystems, which limits the ability of ML models to automatically extract generalizable relationships between features. Second, traditional ML approaches are unable to harness inconsistent feature inputs. Current studies on local environmental ecosystems often enhance the prediction of their target study region by leveraging multiple types of input features related to the target variable. However, the selected features often vary between different studies. Concretely, besides meteorological data (e.g., solar radiation) that are commonly used to drive physics-based models, prior works also explored including other measurements on the environment and target system (e.g., soil properties Chen et al. (2023), land use and geometric structures Cho and Kim (2022)), proxy observations derived from remote sensing Mkhabela et al. (2011), and other physical variables simulated by physics-based models Jia et al. (2021b). These features may not always be available for every data sample, which poses a challenge in training a global model that can utilize different input features. Finally, there is a lack of a general pipeline for dynamically incorporating auxiliary observations into the current methods, hindering their ability to achieve better performance in long-term prediction. Current works on assimilating auxiliary observations rely on task-specific learning mechanisms or model structures, e.g., Kalman filtering for incorporating new observations Zwart et al. (2023), and graph convolution and invertible network layers for assimilating observations from neighboring samples Chen et al. (2021a). Yet these approaches can be computationally expensive to implement when we have a large state space and/or non-linear system dynamics.
In this paper, we propose a novel method, Foundational semantic Recognition for modeling Environmental Ecosystems (FREE) to address these limitations. The key idea of FREE is to translate the heterogeneous input data into natural language descriptions using large language models (LLMs) and then estimate the target variable through semantic recognition in the text space. This enables addressing the challenges in utilizing different data sources by only manipulating the text space while maintaining the same predictive modeling component (i.e., semantic recognition). In particular, the translation process uses the available features of each data point with allowing different or incomplete input features for different data points. Moreover, it can easily incorporate auxiliary observations (e.g., newly collected observations from the previous day) into the textual description with a properly designed prompt. For semantic recognition, we utilize a separate language model to embed the obtained textual description and then create additional network layers (e.g., long-short term memory (LSTM)) to capture data dependencies. The use of the language model on textual descriptions enables a better capture of the nature and semantics of input features. To further enhance the embedding performance of the language model on environmental descriptions, we pre-train the semantic recognition component using abundant simulated samples generated by physics-based models. This pre-training process also helps the model better learn the general physical relationships encoded in the physics-based models and mitigate the challenge posed by the sparse observations.
We evaluate the proposed method on two real-world datasets: (1) stream water temperature data from the Delaware River Basin, and (2) corn yield data in Illinois and Iowa. Both of them cover a diverse set of locations and long time periods. The results demonstrate the superior predictive performance of the proposed FREE method over multiple baselines, especially with sparse observations. The method is shown to be able to effectively handle different input features and incorporate auxiliary observations. The pre-training process also helps improve the prediction performance of the model when adapted to different locations.
2 Related Work
Physics-based models have been widely used to simulate complex environmental processes, including different components in water cycles and plant growth in agroecosystems Zhou et al. (2021); Srinivasan et al. (2010). However, these models are necessarily approximations of reality and often rely on approximations and parameterizations Gupta and Nearing (2014). Recent advancements in ML offer new opportunities for modeling environmental ecosystems using the rapidly growing earth observation data and ground observations about water, plants, soils, and climate. In particular, prior research has shown the promise of ML-based approaches in modeling agroecosystems Jia et al. (2019); Liu et al. (2022); Ghazaryan et al. (2020) and freshwater ecosystems Rahmani et al. (2021).
Data assimilation has been widely used for assimilating auxiliary observations to optimally adjust the model state of a system. There are many data assimilation approaches used for modeling environmental ecosystems, and they differ in their assumptions about the distributions of the model, data predictions, and speed of computation. The Kalman filter, which iteratively estimates the next state through a forward process and then updates the state using new observations, is a popular data assimilation technique. Many variations of the Kalman filter have also been implemented Evensen and others (2009). Despite their extensive use, these approaches can be computationally expensive when we have a large state space and/or non-linear system dynamics. Recently, researchers have started to use neural networks as an alternative way for implementing data assimilation Brajard et al. (2020); Chen et al. (2021b). Note that this is different from online incremental learning in that model parameters remain the same but only states are changed. The intuition of data assimilation is to adjust the model state when they are disturbed by external factors that are not captured by input features.
Figure 1: Workflow of FREE. Input features are first transformed into natural language descriptions. These descriptions are then processed by a separate language model (LM) to generate embeddings, which are fed to an LSTM layer for making predictions. Simulated data generated by a physics-based model are used to pre-train the LM and LSTM layers, followed by fine-tuning with true observations of the target variable.
To mitigate the need for large training data in ML models, one promising direction is to transfer knowledge embodied in physical simulations to ML models. Simulated data can be generated by physics-based models that represent underlying physical processes using a series of mathematical equations. For example, prior work has shown that ML models pre-trained using simulated data can perform much better under data-scarce scenarios for modeling freshwater and agricultural ecosystems Jia et al. (2021a); Read et al. (2019).
Recent advancements in LLMs have demonstrated remarkable performance in solving language tasks based on human instructions Touvron et al. (2023); Chiang et al. (2023) and have been applied to geoscience applications Mai et al. (2023). More recently, these models have been applied to other types of data. For example, LLMs have been used to textualize tabular information and handle question-answering tasks Borisov et al. (2022); Zhao et al. (2023). Textualized tabular data offers the advantage of being able to handle inputs with different feature sets and is more robust in dealing with missing values.
3 Method
In this section, we define the problem and introduce our proposed framework. The objective is to predict the target variable $y$ (e.g., water temperature or crop yield) at multiple locations $i\in\{1,...,N\}$ and over a period of time. For clarity, we denote by $\mathbf{x}_{i,t}$ the input features at the location $i$ on date $t$. The input features contain the set of meteorological variables (e.g., solar radiation, rainfall) that are commonly used as drivers for physics-based models. Besides, we include some other physical variables estimated by physics-based models (e.g., cloud cover fraction, groundwater properties) or obtained from other data sources (e.g., soil properties). Notably, these features may be absent for certain locations or time steps. The observations of the target variable at each location $i$ over multiple time steps is $y_{i}=\{{y}_{i,1},{y}_{i,2},...,{y}_{i,T}\}$ which are only sparsely available for certain time steps and locations.
As outlined in Fig. 1, the proposed framework FREE converts the traditional predictive modeling task in environmental science to the semantic recognition problem in a text space created by LLMs (e.g., GPT). A clear benefit of this framework is its ability to harness inputs of different feature sets and incorporate auxiliary information, e.g., new observations collected from the previous day. We also introduce additional model components (e.g., BERT and LSTM) to embed the text data and capture the temporal dependencies. To mitigate the need for large training samples for different ecosystems, we propose to pre-train the model using abundant simulated data generated by existing physics-based models. In the following, we will describe these components in detail.
Figure 2: Input data conversion. Green arrows indicate FREE handling inputs of diverse feature sets with three major inputs in the prompt: prefix $p$, linearized data, and suffix $s$. Red arrow suggests that traditional ML models need separate preprocessing methodologies (e.g., imputation) to address data irregularities effectively.
3.1 Overall architecture
The proposed FREE framework (shown in Fig. 1) consists of two major components in its architecture.
Input data conversion: To address inconsistencies in the feature set and incorporate auxiliary observations, we propose to transform the original data sample into a corresponding natural language description. This approach facilitates the handling of diverse and potentially incomplete feature sets for different data points, enabling a uniform textual representation of data across varying input scenarios.
Specifically, we leverage existing LLM GPT-3.5 to transform each data point, represented as $\mathbf{x}_{i,t}$, into clear, natural language descriptions $d_{i,t}$. To effectively communicate with GPT-3.5, we use a linearization technique Wang et al. (2023) to construct prompts that consist of a context-setting prefix, the linearized data input, and a directive suffix. The prefix ($p$) provides the model with a background of the dataset, while the suffix ($s$) instructs the model on how to format its output. The complete prompt is thus formulated as:
$d_{i,t}=\text{LLM}\left(p,\text{linearize}(\mathbf{x}_{i,t}),s\right).$ (1)
As depicted in Fig. 2, we format the input features $\mathbf{x}_{i,t}$ of $K$ dimensions into a sequence pairing column names ${c}^{k}_{i,t}$ with the corresponding feature values $\mathbf{x}^{k}_{i,t}$ (e.g., { [rainfall: 0], [solar radiation: 151.14154] } ). This process can be expressed as follows:
$\text{linearize}(\mathbf{x}_{i,t})=\{[{c}^{k}_{i,t}:\mathbf{x}^{k}_{i,t}]\}_{k=1}^{K}.$ (2)
Given the prompt, the LLM can generate feature summaries that are both descriptive and succinct, such as: ”On December 4, 2006, there was no recorded rainfall in the Delaware River Basin. The average air temperature was -3.364 degrees Celsius, where freezing-thawing and phase change may occur. The solar radiation measured was 108.26 watts per square meter.”.
The obtained textual descriptions, alongside the observations in the original dataset, form paired data samples as $\{d_{i,t},y_{i,t}\}$, which are then used for tuning the semantic recognition component (to be discussed later).
In summary, the proposed input conversion process helps generate an understandable and flexible representation of input features. More importantly, the use of LLM enables supplementing the raw feature values with semantically meaningful text descriptions and interpreting underlying physical phenomena, which facilitates subsequent modeling components to capture the complex nature and the interactions of input features.
Semantic recognition: Once obtaining the textual descriptions $d_{i,t}$, we utilize a separate language model (LM) $f$ to process the obtained text data and embed them as
$\mathbf{e}_{i,t}=f(d_{i,t}).$ (3)
In this work, we use the DistilBERT model Sanh et al. (2019) in this step. It is vital to note that the embeddings $\mathbf{e}_{i,t}$ are generated independently across different data points. To effectively capture the intricate data dependencies, we introduce additional network layers that combine multiple data embeddings. In particular, this work uses LSTM layers to capture common temporal data dependencies in the environmental ecosystem, e.g., seasonal changes, and the effect of heavy rainfall on soil water in the next few days. We will explore other advanced network structures, e.g., graph neural networks for modeling streams Jia et al. (2021b) and crop yield Fan et al. (2022), as future work.
3.2 Handling diverse inputs and auxiliary data
One major goal of modeling environmental ecosystems is long-term prediction in large regions. There are opportunities to enhance ML models towards this goal from two aspects. First, predictive models often have larger predictive errors as time progresses in the test phase due to the accumulated bias in the model state. One could explore leveraging newly collected observations during the testing phase (e.g., the observations collected at time $t$) to mitigate the current model bias and improve the future prediction (e.g., the prediction at $t+1$). Second, for some data samples, we may have access to additional features such as local soil properties or modeled variables from physics-based models. These additional features can be combined with the original features to improve the model prediction. Under the proposed FREE framework, we can make use of the additional observations and features by simply modifying the prompt while keeping the predictive model (i.e., the semantic recognition component) the same.
3.2.1 Incorporating auxiliary observations
In this study, we consider incorporating two types of auxiliary observations to enhance the prediction of the data sample $\mathbf{x}_{i,t}$: (1) the newly collected observation from the previous time $t-1$ and at the same location $i$, and (2) the newly collected observations from the neighboring locations of the target location $i$. We incorporate the information from auxiliary data by updating the linearized data input to include it.
As observations are collected from different dates (e.g., the previous day) with the existing features in $\mathbf{x}_{i,t}$, we need to explicitly include the exact date information for the auxiliary observations and original features. If we consider the auxiliary observations being collected at the date $t-1$ from both the current location $i$ and its neighbor $j\in\mathcal{N}(i)$, the updated linearization process can be expressed as
$\text{linearize}([\mathbf{x}_{i,t},y_{i,t-1},y_{j,t-1}]_{j\in\mathcal{N}(i)})$ (4)
$=$ $[\text{date}:t-1]\cup[{c}_{i}^{y}:y_{i,t-1}]\cup[{c}^{y}_{j}:y_{j,t-1}]$
$\cup[\text{date}:t]\cup\{[{c}^{k}_{i,t}:\mathbf{x}^{k}_{i,t}]\}_{k=1}^{K},$
[…]
It is noteworthy that the auxiliary observations may not always be available. When creating the linearized data, we include the columns of the auxiliary observations only if they are available and skip them otherwise.
3.2.2 Handling different input features
FREE can also handle different input features by using only the available features to create the linearized data while skipping missing features. In particular, if a feature $k$ is not available for the data sample $\mathbf{x}_{i,t}$, then the pair $[c^{k}_{i,t}:\mathbf{x}^{k}_{i,t}]$ will be skipped and not included in the input to the LLM. This entails the LLM generating natural descriptions that focus only on the available features.
The proposed approach enables the use of a combination of heterogeneous data sources for modeling complex environmental ecosystems. It also allows the subsequent semantic recognition to proceed seamlessly on the generated text without the need for manual adjustments to account for data irregularities.
3.3 Pre-training using physical simulations
The pre-trained LMs in the semantic recognition phase (Eq. 3) are not inherently trained on data specific to the target environmental ecosystems. As shown in Fig. 3 (b), the input features from samples in different seasons tend to be mixed in the latent embedding space of the original pre-trained DistilBERT model. This indicates that a standard LM, when applied directly, may fall short of capturing semantics from the text generated for our target task.
Tuning the LM towards the target domain requires sufficient observed data samples, which are often not available in real-world ecosystems. To address this, we turn to the simulated data $\{\mathbf{x}_{i,t},{\tilde{y}}_{i,t}\}$, generated by a physics-based model, and use them to pre-train the semantic recognition modeling component. Fig. 3 (a) shows that samples of different seasons can now be distinguished in the updated embedding space, which confirms that the LM model tuned with simulated data can better capture semantics in the textual descriptions $d_{i,t}$.
This simulation-based pre-training approach also helps in facilitating the training of the semantic recognition component to emulate general physical processes encoded in the physics-based model. This helps enhance the model’s generalizability as many physical processes generally hold across space and time. Upon completing this pre-training, we find that it requires only a few epochs of fine-tuning using true observations before reaching a quality model.
Figure 3: Distribution of data points randomly sampled from summer and winter in their embedding space visualized by t-SNE. (a) demonstrates a pronounced segregation between different seasons after domain-specific tuning. (b) shows an overlapping distribution of embeddings created by the original DistilBERT model.
4 Experiments
In this section, we present our datasets and provide a comprehensive assessment of the proposed methods. Our primary emphasis lies on predictive performance using sparse data, underscoring the efficacy of the proposed FREE framework and the simulation-based pre-training. Besides, we delve into additional experiments that validate our model’s capability to harness diverse input features and auxiliary observations. We further examine its performance under transfer learning scenarios and analyze its performance sensitivity by freezing the LM in the semantic recognition component during the fine-tuning phase.
4.1 Experimental Setup
We first introduce two datasets for evaluation: one for stream water temperature prediction and the other for annual crop yield prediction. Then, we introduce the baseline methods used in this paper and the implementation details.
Stream dataset
The dataset is collected from the Delaware River Basin, an ecologically diverse region and a watershed along the east coast of the United States. The datasets are pulled from the U.S. Geological Survey’s National Water Information System US Geological Survey (2016) and the Water Quality Portal Read et al. (2017), the largest standardized water quality data set for inland and coastal waterbodies Read et al. (2017). Observations at a specific latitude and longitude were matched to river segments that vary in length from 48 to 23,120 meters. These segments were defined by the national geospatial fabric used for the National Hydrologic Model as described by Regan and others (2018), and the river segments are split up to have at maximum a one-day water travel time. We match observations to river segments by snapping observations to the nearest stream segment within a tolerance of 250 meters, excluding those over 5,000 meters from a segment’s outlet. Segments with multiple observation sites were aggregated to a single daily average water temperature value.
We conduct our study on a subset of river segments of the Delaware River Basin, specifically focusing on the river segments that converge into the main stem of the Delaware River at Wilmington, DE. This large dataset is referred to as the Christina River Watershed (CRW), which contains 42 river segments. Separately, we also studied a distinct stream segment, named “single stream (SS)”, which has the largest number of water temperature observations. This allows us to test the capacity of models under data-sparse scenarios by intentionally withholding a fraction of training samples.
Our dataset spans a long period from October 31, 2006, to March 30, 2020 (4900 days). On a daily scale, we incorporated basic meteorological features (the day of the year, rainfall, daily average air temperature, and solar radiation). In addition to these, we also considered other features, namely average cloud cover fraction, groundwater temperature, subsurface temperature, and potential evapotranspiration.
While daily average water temperature observations were present for certain segments, they were not consistently available for every date. The count of temperature readings for each segment varied, ranging from a minimum of 0 to a maximum of 4,682. In totality, “CRW” comprises 205,800 temperature observations distributed across all its segments and dates, and “SS” has 4,676 temperature observations. Besides true observations, we also use the simulated water temperature data for CRW and SS generated by the physics-based PRMS-SNTemp model Markstrom (2012). For the following experiments, the pre-training process used simulated data from October 31, 2006, to March 30, 2020. We train each ML model using data from the first 2,450 days (October 31, 2006 to July 8, 2013) and then test in the next 2,450 days (July 8, 2013 to March 30, 2020).
Stream baselines
We compare model performance to multiple state-of-the-art baselines developed for stream modeling, including the Basic LSTM model (LSTM) (considered to be the state of the art in many environmental modeling problems Shen and Lawson (2021)), TransformerVaswani et al. (2017), Recurrent Graph Model (RGRN) Jia et al. (2021b), Graph-temporal convolutional network (Gr-CNN) Sun et al. (2021), HydroNets Moshe et al. (2020). Besides, we include a baseline method that assimilates new observations for updating the model, the Heterogeneous Recurrent Graph Networks with data assimilation (HRGN-DA) Chen et al. (2021a).
Table 1: Prediction RMSE for stream water temperature using 1%, 2%, 4%, and 100% randomly selected training labels. The best results are bold, while the second best results are underlined.
Dataset Method 100% 4% 2% 1%
SS LSTM 1.90 1.95 2.14 2.23
Transformer 1.79 2.05 2.20 2.27
FREE 1.70 1.71 1.90 2.09
CRW LSTM 1.90 2.33 2.80 2.87
Transformer 2.14 2.20 2.19 2.30
RGRN 1.78 2.31 2.61 2.80
Gr-CNN 1.80 2.38 2.69 3.56
HydroNets 1.87 2.51 2.77 3.75
FREE 1.61 1.68 1.62 1.65
Table 2: Prediction RMSE for annual corn yield using 10%, 20%, 50%, and 100% randomly selected training labels. The best results are bold, while the second best results are underlined.
Dataset Method 100% 50% 20% 10%
Crop LSTM 73.56 79.67 78.88 84.52
PG-GNN 79.92 71.68 72.72 79.04
PG-AN 96.83 114.40 90.52 100.36
Transformer 84.93 79.78 97.56 101.40
FREE 65.24 63.74 63.75 64.28
Crop dataset
We use the corn yield data in Illinois and Iowa from the years 2000-2020 provided by USDA National Agricultural Statistics Service (NASS)  [^1] https://quickstats.nass.usda.gov/. In particular, there are in total 199 counties in our study region (100 counties in Illinois and 99 counties in Iowa). The corn yield data (in gCm-2) are available for each county each year. The observed annual crop yield labels cover 199 countries spanning 21 years from 2000-2020, with daily records for each year. The input data include seven features, namely, the surface downward shortwave radiation, the maximum air temperature, the maximum humidity, the wind speed, the precipitation, the depth-weighted averaged bulk density in the soil, and the depth-weighted averaged sand content in the soil. The feature data collected from NLDAS-2 climate data Xia and others (2012), 0-30cm gSSURGO soil properties [^2] https://gdg.sc.egov.usda.gov/, and the crop type information.
Besides the true observed yield labels, we use the physics-based Ecosys model Zhou et al. (2021) to simulate ecosystem autotrophic respiration (Ra), ecosystem heterotrophic respiration (Rh), net ecosystem exchange (NEE), and crop yield for 10,335 synthetic sample locations in the United States from the years 2000-2020. All synthetic data are used for pre-training. To reduce computational load, we randomly sample 1/7 of synthetic data in the pre-training phase. We use the observational data from 2000-2017 for training and the data from 2018-2020 for testing.
Crop baselines
We compare the proposed method with the LSTM-based model and two other baselines that have both proven effective in crop yield prediction, i.e., the physics-guided attention network (PG-AN) He et al. (2023), and the graph neural network-based model Fan et al. (2022) pre-trained with the same simulated data (PG-GNN).
Implementation details
We pre-train the FREE model with 800 epochs and fine-tune it with an additional 50 epochs. We implemented our methods using PyTorch == 2.0.1, on an RTX 4090 GPU and employed an ADAM optimizer with a 1e-5 learning rate for DistilBERT, and 0.00025/0.0001 pre-training/fine-tuning learning rates for LSTM. Additionally, the batch size is 1 and the hidden dimension is 256.
4.2 Overall predictive performance
Table 1 and Table 2 show the performance of various methods for predicting stream water temperature and predicting annual crop yield, respectively. The notation FREE represents a model pre-trained on synthetic data and then fine-tuned on observational data. We also conduct the sparsity test on both datasets. For this sparsity test, training was conducted using only {4%, 2%, 1%} and {50%, 20%, 10%} of training data randomly sampled from the whole training period on the stream dataset and the crop dataset, respectively.
According to Table 1, the FREE outperforms all baseline methods when data are abundant, which shows the effectiveness of the proposed method in capturing the semantics and relationships of input features from textual descriptions. Notably, the FREE maintains good performance regardless of label volume, while other models generally have much worse performance in data-sparse scenarios. This underscores the model enhancement brought by the simulation-based pre-training in learning generalizable and domain-specific semantics. Moreover, pre-training our model before fine-tuning significantly reduces the time for effective fine-tuning (about one-tenth of the time compared to direct fine-tuning). Table 2 presents similar comparisons between the proposed method and the baselines in predicting annual corn yield. We also observe that some methods can have slightly better performance with reduced training data. This is because the data samples are highly heterogeneous across different counties and years in this dataset, and the randomly sampled training data have a chance to get slightly closer to the testing data in distribution.
Figure 4: Comparison of pre-trained FREE models and the baseline LSTM on stream water temperature prediction. (a) SS test case with full training data. (b) SS test case with 1% training data. (c) an example segment in CRW with full training data from CRW. (d) the same example segment in CRW with 1% training data from CRW.
Figure 5: Evaluation of FREE with auxiliary information. the variation in test errors for (a) and (b)  predicting stream water temperature upon the inclusion of extra observations under different sparsity levels in SS and CRW separately, (c) predicting stream water temperature upon the integration of additional features, (d) predicting corn yield upon the integration of additional features.
In Fig. 4, we show the predicted water temperature over one year for the SS test case and another river segment within the CRW test case to demonstrate the alignment of predictions and true observations. The figure clearly shows that our method outperforms the baseline model under different data sparsity scenarios (100% and 1%), especially in capturing the water temperature oscillations in the summertime. The LSTM model’s predictions, while reasonably tracking the general trend, fall short in replicating the nuanced variations, especially as the training data becomes sparser. In scenarios with limited data, the LSTM struggles to recognize subtle changes and creates smooth predictions, overlooking smaller peaks and valleys that our method consistently identifies and traces.
4.3 Handling auxiliary information
We also evaluate the performance of our method in harnessing auxiliary observations and diverse input features. FREE-C refers to the model that incorporates auxiliary observed water temperatures of the current river segment from the prior day (when they are available). As shown in Fig. 5 (a) and (b), FREE-C outperforms FREE, demonstrating that our proposed approach can process auxiliary observation data effectively. It is worth mentioning that in Fig. 5 (b), the baseline HRGN-DA, which also incorporates new observational data, initially achieves close performance with the proposed FREE-C approach using the complete training dataset. Nonetheless, when the available labeled data is reduced, our FREE-C models surpass HRGN-DA in accuracy. Additionally, HRGN-DA is designed for application to a graph of streams, making it incompatible with the SS dataset. For crop yield prediction, unlike daily stream temperature prediction, crop yield labels are available on a yearly scale. The yield labels from the previous year usually have little impact on the yield of the next year and thus we did not include crop yield’s auxiliary observations experiments.
[…]
Referencing Fig. 5, two primary observations are evident: (1) The integration of auxiliary observations improves the performance of the stream modeling dataset. Notably, FREE-C achieves the best results overall in both the CRW and SS datasets. (2) For both datasets, the use of additional features can reduce the prediction errors. We will explore using more diverse input features in future work.
Figure 6: Comparison between FREE, the non-pre-trained FREE, and the transferred FREE (pre-trained from CRW) under different sparsity levels in the target SS. (a) train on meteorological features. (b) trained on meteorological and four additional features.
4.4 Evaluation on model transferability
We assess the performance of adapting a model pre-trained using simulated data from CRW to the river segment in a different region (SS). In particular, we compared the transferred models (FREEtrs and FREE-A4trs) to the models trained on observational data directly without pre-training (FREEnprt and FREE-A4nprt) and the models pre-trained on the same segment in SS (FREE, FREE-A4). Here we use the subscript “trs” to represent the model being pre-trained on CRW. As Fig. 6 shows, The pre-trained models (FREEtrs, FREE-A4trs, FREE, FREE-A4) consistently outperform the models FREEnprt, FREE-A4nprt under different data sparsity levels, maintaining stable performance despite reduced label availability. Remarkably, the pre-trained model from the different domain CRW can have a similar degree of contribution to performance enhancement compared to the model pre-trained on the same target segment. These findings emphasize the pre-trained models’ ability to not only address the challenges caused by sparse labels but also to transfer learning capabilities. This shows the potential of building a global pre-trained model over large regions, which can then be adapted efficiently to different target regions with small training data.
4.5 Fine-tuning analysis
We investigate the effect of two different fine-tuning strategies: (1) fine-tuning all the parameters of the semantic recognition model, i.e., LM+LSTM (FT-All), and (2) fine-tuning only the parameters of the LSTM layer while fixing the embeddings created by the LM model (FT-LSTM). The underlying assumption for the fine-tuning approach (2) is that the LM can generate semantics-aware embeddings after being pre-trained on simulations, eliminating the need to be further adjusted on true observations. The results presented in Table 3 show that comprehensive fine-tuning FT-All over all the parameters within our pre-trained models yields better results. Interestingly, the partial fine-tuning approach FT-LSTM, where we fix the parameters of the LM and only adjust those of the LSTM, also delivers decent performance. While this tuning approach (FT-LSTM) slightly compromises the predictive accuracy compared to the comprehensive fine-tuning approach (FT-All), it requires much less computational cost, making it a convenient and practical alternative in resource-limited scenarios.
Table 3: Prediction RMSE for stream water temperature and crop yield using different fine-tuned models. FT-All means fine-tuning the whole model using training labels, and FT-LSTM means only fine-tuning the LSTM layer.
Dataset Method FT-All FT-LSTM
SS FREE 1.70 1.71
FREE-C 1.11 1.70
FREE-A2 1.73 1.98
FREE-A3 1.68 1.93
FREE-A4 1.63 1.94
CRW FREE 1.61 2.22
FREE-C 1.02 2.35
FREE-U 1.50 2.28
FREE-CU 1.03 2.27
FREE-A2 1.68 2.18
FREE-A3 1.61 1.98
FREE-A4 1.65 2.11
Crop FREE-A0 65.70 66.73
FREE-A1 65.45 66.68
FREE-A2 65.24 66.01
5 Conclusion
This paper introduces FREE, a novel LLM-based framework for modeling environmental ecosystems. FREE is shown to outperform existing methods in the context of predicting stream water temperature and annual crop yield. This demonstrates the potential of FREE in better capturing the semantics of environmental features by converting them to natural language descriptions. FREE also has the flexibility to utilize different input features and auxiliary observations for different data samples to enhance predictive performance. Moreover, FREE maintains high predictive performance even under data-sparse scenarios. This is attributed to the model initialization through the pre-training on simulated data. The simulation-based pre-training also aids in extracting physically consistent data patterns, which improves the generalizability and transferability to different regions.
We anticipate the FREE framework to serve as a stepping stone to building foundational solutions for modeling complex environmental and physical systems. This framework remains limited in (1) its computational cost (e.g., caused by the increasing length of textual description) when combining a large number of physical variables as inputs, (2) the ability to integrate known physical and causal relationships in the complex ecosystems, and (3) effectively handling highly erratic or noisy data beyond incomplete or missing data scenarios. We will investigate these directions in future work.